OpenStudy (anonymous):

solve this equation 8 = -4 +x

OpenStudy (anonymous):

Do you think you can isolate the variable? ( Get x by itself? )

OpenStudy (mathstudent55):

Since the x variable is on the right side, first switch sides. Then notice that -4 is being added to x. What do you add to -4 to get zero? That's what you must add to both sides to get rid of the -4.

OpenStudy (anonymous):

X=12

OpenStudy (anonymous):

^ x - 4 = 8 x = 8 + 4 x = 12